# DeepEdit: Knowledge Editing as Decoding with Constraints

### v2

@@ 5.5 Case Study
Figure 8 gives a qualitative comparison example between MeLLo and our DeepEdit on MQuaKE-3k. The results show that the strong KE method MeLLo makes mistakes on KE. For example, given the question “The writer of the novel “1984” died in which continent? ”, MeLLo is unable to generate the helpful sub-question to find the final correct answer. In contrast, our DeepEdit, based on the constraints of [Conciseness], [Coherence], and [Receptiveness], correctly find the relevant new knowledge “Bucharest is located in the continent of North America.” and obtain the correct answer North America with respect to the new knowledge.
@@
Taking a closer look, we observe that our DeepEdit’s induced reasoning is much more succinct than MeLLo, since we effectively filter out the irrelevant and redundant information during decoding. Because longer and noisy textual information can lead to more serious hallucinations and higher difficulty of reasoning, our DeepEdit exhibits a stronger capability of KE for LLMs. Overall, our DeepEdit method enforces LLMs’ outputs to follow our decoding constraints so as to soundly incorporate new knowledge in reasoning, which leads to the correct predictions in Figure 8.
6 Conclusion
We explore a new paradigm for knowledge editing of black-box LLMs: decoding with constraints. We have designed a new decoding method DeepEdit to improve the knowledge editing of LLMs. We enforce the LLM’s outputs to satisfy the desired properties with decoding at the reasoning step level. Based on the DeepEdit, we can improve the LLMs’ reasoning with new knowledge though higher conciseness, coherence, pertinence, and receptiveness. Our DeepEdit controls LLMs’ reasoning to soundly incorporate new knowledge by satisfying our constraints. In addition to DeepEdit, we provide two new benchmarks: MQuAKE-2002 and MQuAKE-hard to provide more precise and challenging evaluation for KE methods respectively. Extensive experiments demonstrate the effectiveness of our DeepEdit. Future work includes analyzing the effects of other constraints that can improve knowledge editing for large language models.

### v3

@@ Appendix D Case Study
Figure 9 gives a qualitative comparison example between MeLLo and our DeepEdit on MQuaKE-3k. The results show that the strong KE method MeLLo makes mistakes on KE. For example, given the question “The writer of the novel “1984” died in which continent? ”, MeLLo is unable to generate the helpful sub-question to find the final correct answer. In contrast, our DeepEdit, effectively controls LLMs to correctly find the relevant new knowledge “Bucharest is located in the continent of North America.” and obtain the correct answer North America with respect to the new knowledge.
@@
Taking a closer look, we observe that our DeepEdit’s induced reasoning is much more succinct than MeLLo, since we effectively filter out the irrelevant and redundant information during decoding. Because longer and noisy textual information can lead to more serious hallucinations and higher difficulty of reasoning, our DeepEdit exhibits a more effective KE than MeLLo with its succinct reasoning process controlled by constraints. Overall, our DeepEdit enhances LLMs to follow our decoding constraints so as to soundly incorporate new knowledge in multi-step reasoning, which leads to the correct predictions in Figure 9.